OpenStudy (anonymous):

A group of 8 friends (5 girls and 3 boys) plans to watch a movie, but they have only 5 tickets. If they randomly decide who will watch the movie, what is the probability that there are at least 3 girls in the group that watch the movie?

OpenStudy (anonymous):

A group of 8 friends (5 girls and 3 boys) plans to watch a movie, but they have only 5 tickets. If they randomly decide who will watch the movie, what is the probability that there are at least 3 girls in the group that watch the movie? ------ # of ways to succeed:: 5C3*3C2 = 5C2*3C2 = 10*3 = 30 ways ---- # of random groups of 5: 8C5 = 8C3 = (8*7*6)/(1*2*3) = 56 ====== Ans: P(3 girls and 2 boys) = 30/56 = 0.5357